# Magneto-optical transport in tilted type-I double-Weyl semimetals in the presence of orbital magnetic moment

### v1

@@ IV Second order non-linear response of DWSMs
$\sigma_{zxz}^{(B)}(2\omega)=s\frac{2\mathit{v}_{F}^{2}}{3\pi\mu\alpha_{2}}\sigma_{2}^{(B)}\cos\gamma$ (61)
$\sigma_{xzz}^{(B)}(2\omega)=-s\frac{4\mathit{v}_{F}^{2}}{3\pi\mu\alpha_{2}}\sigma_{2}^{(B)}\cos\gamma$ (62)
$\sigma_{zyz}^{(B)}(2\omega)=s\frac{2\mathit{v}_{F}^{2}}{3\pi\mu\alpha_{2}}\sigma_{2}^{(B)}\sin\gamma$ (63)
$\sigma_{yzz}^{(B)}(2\omega)=-s\frac{8\mathit{v}_{F}^{2}}{3\pi\mu\alpha_{2}}\sigma_{2}^{(B)}\sin\gamma$ (64)
In this case, the B-linear contribution to the second harmonic conductivity is dependent of the chirality but independent of the tilt. Summing the conductivity over the Weyl cones with opposite chirality cancels each other, leading to the disappearance of the total B-linear contribution to the second harmonic conductivity.
IV.3 The second order nonlinear Hall conductivity $\sigma_{abc}^{(H,0)}$ in the absence of magnetc field
In this subsection, we study the second-order nonlinear Hall effect of Weyl semimetals wthout magnetic field. Inserting Eq. (III) into the second term in Eq. (IV) and taking B = 0, we obtain the nonlinear Hall conductivity.
@@ IV.3 The second order nonlinear Hall conductivity $\sigma_{abc}^{(H,0)}$ in the absence of magnetc field
where $D_{bd}$ is called Berry dipole. It is a first order moment of the Berry curvature[31].
@@
In the case of DWSMs,the magntude of Berry curvature dipole components are twice the values of single WSMs and it has the same dependence on the chirality, as well as on the tilt. Therefore, contributions from a pair of Weyl nodes with opposite chirality exactly cancel each other [22, 31, 32].
@@ IV.4 Linear B-contribution to the second order nonlinear Hall conductivity $\sigma_{abc}^{(H,B)}$
Now, we focus on the second-order nonlinear Hall effect in a weak magnetic field. Inserting Eq. (III) into the second term of Eq. (IV) and retain terms up to first power of B, one obtan complex nonlinear Hall conductivity
@@
where $D_{bd}^{\Omega}$ and $D_{bd}^{m}$ are Berry curvature dipole contributions due to the Berry-curvature and the orbital magnetic moment respectively.
@@
Case-I In the case of $\textbf{B}\parallel t_{s}$ , one obtain the Berry curvature dipole components
$D_{xx}^{\Omega}=D_{yy}^{\Omega}=D_{zz}^{\Omega}=0$ (72)
@@
In this case all the components are zero which are finite in case of isotropic WSM[22].
@@
$D_{zx}^{\Omega}$ $=$ $-\frac{4}{3}t_{s}D_{2}^{(B)}\cos\gamma$ (74)
$D_{xz}^{\Omega}$ $=$ $\frac{15\pi}{2}\frac{\mu\alpha_{2}}{\mathit{v}_{F}^{2}\hbar}t_{s}D_{2}^{(B)}\cos\gamma$ (75)
$D_{zy}^{\Omega}$ $=$ $-\frac{4}{3}t_{s}D_{2}^{(B)}\sin\gamma$ (76)
$D_{yz}^{\Omega}$ $=$ $-\frac{15\pi}{2}\frac{\mu\alpha_{2}}{\mathit{v}_{F}^{2}\hbar}t_{s}D_{2}^{(B)}\sin\gamma$ (77)
@@
$D_{zx}^{m}=-6t_{s}D_{2}^{(B)}\cos\gamma$ (78)
$D_{zy}^{m}=-6t_{s}D_{2}^{(B)}\sin\gamma$ (79)
where $D_{2}^{(B)}=\frac{1}{8\pi^{2}}\frac{eB}{\hbar}\frac{\hbar^{2}\mathit{v}_{F}^{2}}{15\mu^{2}}$. One can check that $D_{zx}^{m}=D_{xz}^{m}$ and $D_{zy}^{m}=D_{yz}^{m}$ and rest of the components will vanish. The contributions of all components are renormalized due to parameters $\alpha_{2}$ and $\mathit{v}_{F}$ of DWSMs as compared to single WSMs [22].
@@
Assume that an electromagnetic wave propogates in the x direction:
@@
is the Jones vector in the y-z plane with phases $\alpha_{y},\alpha_{z}$ , and the amplitudes $E_{y}=|\textbf{E}|\sin\theta$ and $E_{z}=|\textbf{E}|\cos\theta$ Inserting Eq. (81) into Eq. (80), we obtain the nonlinear Hall current
@@
Following the same procedure of Ref.([22]), Eq.(83) can be rewritten as $j_{x}\sim(D_{yy}-D_{zz})E_{y}E_{z}$ Hence conductivity will be
@@
In contrast to the case of single WSM, we get zero value of conductivity.

### v2

@@ IV.2 Second harmonic conductivity $\sigma_{abc}^{B}$ in the presence of linear magnetic field
for J=1
$\sigma_{zxz}^{(B)}(2\omega)$ $=$ $s\frac{\mathit{v}_{F}^{2}}{12\pi^{2}\mu}\sigma_{2}^{(B)}\cos\gamma$ (110)
$\sigma_{xzz}^{(B)}(2\omega)$ $=$ $-s\frac{\mathit{v}_{F}^{2}}{12\pi^{2}\mu}\sigma_{2}^{(B)}\cos\gamma$ (111)
$\sigma_{zyz}^{(B)}(2\omega)$ $=$ $s\frac{\mathit{v}_{F}^{2}}{12\pi^{2}\mu}\sigma_{2}^{(B)}\sin\gamma$ (112)
$\sigma_{yzz}^{(B)}(2\omega)$ $=$ $-s\frac{\mathit{v}_{F}^{2}}{12\pi^{2}\mu}\sigma_{2}^{(B)}\sin\gamma$ (113)
for J=2
$\sigma_{zxz}^{(B)}(2\omega)$ $=$ $s\frac{\mathit{v}_{F}^{2}}{6\pi^{2}\mu}\sigma_{2}^{(B)}\cos\gamma$ (114)
$\sigma_{xzz}^{(B)}(2\omega)$ $=$ $-s\frac{2\mathit{v}_{F}^{2}}{3\pi^{2}\mu}\sigma_{2}^{(B)}\cos\gamma$ (115)
$\sigma_{zyz}^{(B)}(2\omega)$ $=$ $s\frac{\mathit{v}_{F}^{2}}{6\pi^{2}\mu}\sigma_{2}^{(B)}\sin\gamma$ (116)
$\sigma_{yzz}^{(B)}(2\omega)$ $=$ $-s\frac{2\mathit{v}_{F}^{2}}{3\pi^{2}\mu}\sigma_{2}^{(B)}\sin\gamma$ (117)
for J=3
$\sigma_{zxz}^{(B)}(2\omega)$ $=$ $s\frac{\mathit{v}_{F}^{2}}{4\pi^{2}\mu}\sigma_{2}^{(B)}\cos\gamma$ (118)
$\sigma_{xzz}^{(B)}(2\omega)$ $=$ $-s\frac{\mathit{v}_{F}^{2}}{4\pi^{2}\mu}\sigma_{2}^{(B)}\cos\gamma$ (119)
$\sigma_{zyz}^{(B)}(2\omega)$ $=$ $s\frac{\mathit{v}_{F}^{2}}{4\pi^{2}\mu}\sigma_{2}^{(B)}\sin\gamma$ (120)
$\sigma_{xzz}^{(B)}(2\omega)$ $=$ $-s\frac{\mathit{v}_{F}^{2}}{4\pi^{2}\mu}\sigma_{2}^{(B)}\sin\gamma$ (121)
So, we can rewrite the above components follow the power laws $\sigma_{zxz}^{(B)}(2\omega)\propto\frac{J}{\mu}$ , $\sigma_{xzz}^{(B)}(2\omega)\propto\frac{J}{\mu}$ , $\sigma_{zyz}^{(B)}(2\omega)\propto\frac{J}{\mu}$ and $\sigma_{yzz}^{(B)}(2\omega)\propto\frac{J}{\mu}$.
In this case, the B-linear contribution to the second harmonic conductivity is dependent on the chirality but independent of the tilt. Summing the conductivity over the Weyl cones with opposite chirality cancels each other, leading to the disappearance of the total B-linear contribution to the second harmonic conductivity.
IV.3 The second order nonlinear Hall conductivity $\sigma_{abc}^{(H,0)}$ in the absence of magnetic field
In this subsection, we study the second-order nonlinear Hall effect of Weyl semimetals without magnetic field. Inserting Eq. (16) into the second term in Eq. (94) and taking B = 0, we obtain the nonlinear Hall conductivity.
@@ IV.3 The second order nonlinear Hall conductivity $\sigma_{abc}^{(H,0)}$ in the absence of magnetic field
where $D_{bd}$ is called Berry dipole. It is a first-order moment of the Berry curvature[40].
@@
for J=1
$D_{xx}=D_{yy}$ $=$ $-s\frac{1}{16\pi^{2}}\Bigl{[}\frac{2}{t_{s}}+\frac{1-t_{s}^{2}}{t_{s}^{3}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$ (124)
$D_{zz}$ $=$ $-s\frac{1}{8\pi^{2}}\frac{t_{s}^{2}-1}{t_{s}^{3}}\Bigl{[}2t_{s}+\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$ (125)
for J=2
@@
for J= 3
$D_{xx}=D_{yy}$ $=$ $-s\frac{1}{8\pi^{2}}\Bigl{[}\frac{3}{t_{s}^{2}}+\frac{3(1-t_{s}^{2})}{2t_{s}^{3}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$ (128)
$D_{zz}$ $=$ $-s\frac{1}{4\pi^{2}}\frac{t_{s}^{2}-1}{t_{s}^{3}}\Bigl{[}3t_{s}+\frac{3}{2}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$ (129)
In the case of DWSMs and Triple-WSMs, the magnitude of Berry curvature dipole components is twice and thrice respectively the values of single WSMs and it has the same dependence on the chirality, as well as on the tilt. Therefore, contributions from a pair of Weyl nodes with opposite chirality exactly cancel each other [22, 40, 41].
@@ IV.4 Linear B-contribution to the second order nonlinear Hall conductivity $\sigma_{abc}^{(H,B)}$
Now, we focus on the second-order nonlinear Hall effect in a weak magnetic field. Inserting Eq. (16) into the second term of Eq. (94) and retaining terms up to the first power of B, one obtains complex nonlinear Hall conductivity
@@
where $D_{bd}^{\Omega}$ and $D_{bd}^{m}$ are Berry curvature dipole contributions due to the Berry curvature and the orbital magnetic moment, respectively.
@@
Case I In the case of $\textbf{B}\parallel t_{s}$, one obtains the Berry curvature dipole components.
for J=1
$D_{xx}^{\Omega}=D_{yy}^{\Omega}$ $=$ $-\frac{t_{s}}{120}D_{2}^{(B)}$ (133)
$D_{zz}^{m}$ $=$ $\frac{t_{s}}{60}D_{2}^{(B)}$ (134)
$D_{xx}^{m}=D_{yy}^{m}$ $=$ $\frac{t_{s}}{60}D_{2}^{(B)}$ (135)
$D_{zz}^{m}$ $=$ $-\frac{t_{s}}{30}D_{2}^{(B)}$ (136)
for J=2
$D_{xx}^{\Omega}=D_{yy}^{\Omega}=D_{zz}^{\Omega}=0$
@@
In this case, all components are zero, which is finite in the case of isotropic WSM[22].
for J=3
$D_{xx}^{\Omega}=D_{yy}^{\Omega}=\frac{D_{2}^{(B)}}{8}\frac{\alpha_{3}^{\frac{2}{3}}\pi^{3/2}\mu^{\frac{4}{3}}}{\hbar^{\frac{4}{3}}\mathit{v_{F}^{2}}}\Bigl{[}\frac{\pi^{\frac{5}{2}}2766400(\frac{1}{1-t_{s}})^{\frac{1}{3}}((-1+t_{s}){{}_{2}F_{1}{(\frac{-2}{3},\frac{8}{3},\frac{16}{3},\frac{2t_{s}}{-1+t_{s}})}}+{{}_{2}F_{1}{(\frac{1}{3},\frac{8}{3},\frac{16}{3},\frac{2t_{s}}{-1+t_{s}})}})}{729(3)^{\frac{1}{2}}t_{s}\Gamma{(\frac{25}{6})}\Gamma{(\frac{19}{3})}}\Bigr{]}$ (139)
$D_{zz}^{\Omega}=-2\frac{D_{2}^{(B)}}{8}\frac{\alpha_{3}^{\frac{2}{3}}\pi^{3/2}\mu^{\frac{4}{3}}}{\hbar^{\frac{4}{3}}\mathit{v_{F}^{2}}}\Bigl{[}\frac{\pi^{\frac{5}{2}}2766400(\frac{1}{1-t_{s}})^{\frac{1}{3}}((-1+t_{s}){{}_{2}F_{1}{(\frac{-2}{3},\frac{8}{3},\frac{16}{3},\frac{2t_{s}}{-1+t_{s}})}}+{{}_{2}F_{1}{(\frac{1}{3},\frac{8}{3},\frac{16}{3},\frac{2t_{s}}{-1+t_{s}})}})}{729(3)^{\frac{1}{2}}t_{s}\Gamma{(\frac{25}{6})}\Gamma{(\frac{19}{3})}}\Bigr{]}$ (140)
$D_{xx}^{m}=D_{yy}^{m}=\frac{D_{2}^{(B)}}{8}\frac{\alpha_{3}^{\frac{2}{3}}\pi^{3/2}\mu^{\frac{4}{3}}}{\hbar^{\frac{4}{3}}\mathit{v_{F}^{2}}}\frac{4(2)^{\frac{1}{3}}\pi}{729(1-t_{s})^{\frac{1}{3}}\Gamma{\bigl{(}\frac{7}{3}\bigr{)}}}\Bigl{[}6561\Gamma{\bigl{(}\frac{2}{3}\bigr{)}^{2}{{}_{2}F_{1}{\Bigl{(}\frac{1}{3},\frac{5}{7},\frac{7}{3},\frac{2t_{s}}{-1+t_{s}}\Bigr{)}}}}+\\ \frac{2(3)^{\frac{1}{2}}\pi\Gamma{\bigl{(}\frac{-1}{3}\bigr{)}}\bigg{\{}(65+3t_{s}(26+3t_{s}(1+72t_{s}))){{}_{2}F_{1}{\Bigl{(}\frac{1}{3},\frac{8}{3},\frac{10}{3},\frac{2t_{s}}{-1+t_{s}}\Bigr{)}}}-(1+t_{s})(65+9t_{s}(13+12t_{s})){{}_{2}F_{1}{\Bigl{(}\frac{4}{3},\frac{8}{3},\frac{10}{3},\frac{2t_{s}}{1+t_{s}}\Bigr{)}}}\bigg{\}}}{t_{s}^{3}\Gamma{\bigl{(}\frac{10}{3}\bigr{)}}}\Bigr{]}$ (141)
$D_{zz}^{m}=-\frac{D_{2}^{(B)}}{8}\frac{\alpha_{3}^{\frac{2}{3}}\pi^{3/2}\mu^{\frac{4}{3}}}{\hbar^{\frac{4}{3}}\mathit{v_{F}^{2}}}\frac{4\pi^{2}(2)^{\frac{1}{3}}\Gamma{(\frac{2}{3})}}{81(3)^{\frac{1}{2}}(1-t_{s})^{\frac{1}{3}}t_{s}^{4}\Gamma{(\frac{7}{2})^{2}}}\Bigl{[}(-91+3t_{s}(-13+t_{s}(29+9t_{s}))){{}_{2}F_{1}{\Bigl{(}\frac{1}{3},\frac{5}{3},\frac{7}{3},\frac{2t_{s}}{-1+t_{s}}\Bigr{)}}}\\ -(1+t_{s})(-91-78t_{s}+9t_{s}^{2}){{}_{2}F_{1}{\Bigl{(}\frac{4}{3},\frac{5}{3},\frac{7}{3},\frac{2t_{s}}{-1+t_{s}}\Bigr{)}}}\Bigr{]}$ (142)
where, $D_{2}^{(B)}=\frac{eB\hbar\mathit{v}_{F}^{2}}{\pi^{2}\mu^{2}}$.
Refer to caption  Refer to caption
Figure 13: The nonlinear Hall conductivities for the process of second harmonic generation as a function of the incident photon frequency at tilt $t_{+}=0.5$ , B=1 T and $\gamma=\frac{\pi}{4}$ for (a) single WSM (b)triple WSM.. The other parameters are the same as those of Fig.(1)
@@
for J=1
$D_{zx}^{\Omega}$ $=$ $-\frac{t_{s}}{20}D_{2}^{(B)}\cos\gamma$ (143)
$D_{xz}^{\Omega}$ $=$ $\frac{3t_{s}}{40}D_{2}^{(B)}\cos\gamma$ (144)
$D_{zy}^{\Omega}$ $=$ $-\frac{t_{s}}{20}D_{2}^{(B)}\sin\gamma$ (145)
$D_{yz}^{\Omega}$ $=$ $\frac{3t_{s}}{40}D_{2}^{(B)}\sin\gamma$ (146)
@@
$D_{zx}^{m}$ $=$ $-\frac{t_{s}}{40}D_{2}^{(B)}\cos\gamma$ (147)
$D_{zy}^{m}$ $=$ $-\frac{t_{s}}{40}D_{2}^{(B)}\sin\gamma$ (148)
for J= 2
$D_{zx}^{\Omega}$ $=$ $-2\frac{t_{s}}{20}D_{2}^{(B)}\cos\gamma$ (149)
$D_{xz}^{\Omega}$ $=$ $\frac{\pi\mu\alpha_{2}}{16\hbar v_{F}^{2}}t_{s}D_{2}^{(B)}\cos\gamma$ (150)
$D_{zy}^{\Omega}$ $=$ $-2\frac{t_{s}}{20}D_{2}^{(B)}\sin\gamma$ (151)
$D_{yz}^{\Omega}$ $=$ $\frac{\pi\mu\alpha_{2}}{16\hbar v_{F}^{2}}t_{s}D_{2}^{(B)}\sin\gamma$ (152)
and
$D_{zx}^{m}=-2\frac{t_{s}}{40}D_{2}^{(B)}\cos\gamma$ (153)
$D_{zy}^{m}=-2\frac{t_{s}}{40}D_{2}^{(B)}\sin\gamma$ (154)
for J=3
$D_{zx}^{\Omega}$ $=$ $-\frac{3t_{s}}{20}D_{2}^{(B)}\cos\gamma$ (155)
$D_{zy}^{\Omega}$ $=$ $-\frac{3t_{s}}{20}D_{2}^{(B)}\sin\gamma$ (156)
$D_{zx}^{m}$ $=$ $-\frac{3t_{s}}{40}D_{2}^{(B)}\cos\gamma$ (157)
$D_{zy}^{m}$ $=$ $-\frac{3t_{s}}{40}D_{2}^{(B)}\sin\gamma$ (158)
Refer to caption  Refer to caption
Figure 14: The angle dependence of the nonlinear Hall conductivities at $\omega=5THz$ for (a) single WSM (b) triple WSM.. The other parameters are the same as those of Fig.(1)
$\begin{split}D_{xz}^{\Omega}&=\frac{D_{2}^{(B)}}{8}\frac{\alpha_{3}^{\frac{2}{3}}\pi^{3/2}\mu^{\frac{4}{3}}}{\hbar^{\frac{4}{3}}\mathit{v_{F}^{2}}}\frac{\Gamma\big{(}\frac{-1}{3}\big{)}^{2}}{486(2)^{\frac{2}{3}}t_{s}^{4}(1-t_{s})^{\frac{1}{3}}\Gamma\big{(}\frac{7}{3}\big{)}}\Bigg{[}\big{\{}-91+3t_{s}(-13+t_{s}(43+3t_{s}(5+2t_{s}(1+9t_{s}))))\big{\}}{{}_{2}F_{1}\big{(}\frac{1}{3},\frac{5}{3},\frac{7}{3},\frac{2t_{s}}{-1+t_{s}}\big{)}}\\ &-(1+t_{s})\big{\{}91+3t_{s}\big{(}-26+t_{s}(17+6t_{s}(2+3t_{s}))\big{)}\big{\}}{{}_{2}F_{1}\big{(}\frac{4}{3},\frac{5}{3},\frac{7}{3},\frac{2t_{s}}{-1+t_{s}}\big{)}}\Bigg{]}\cos\gamma\end{split}$ (159)
$\begin{split}D_{yz}^{\Omega}=\frac{D_{2}^{(B)}}{8}\frac{\alpha_{3}^{\frac{2}{3}}\pi^{3/2}\mu^{\frac{4}{3}}}{\hbar^{\frac{4}{3}}\mathit{v_{F}^{2}}}\Bigg{[}\frac{(3)^{\frac{1}{2}}\Gamma{(\frac{2}{3})^{3}}\big{\{}(-13+15t^{2}+54t^{4}){{}_{2}F_{1}{(\frac{1}{3},\frac{5}{3},\frac{10}{3},\frac{2t}{-1+t_{s}})}}-(1+t_{s}){{}_{2}F_{1}{(\frac{4}{3},\frac{5}{3},\frac{10}{3},\frac{2t_{s}}{-1+t})}}\big{\}}}{{14(2)^{\frac{2}{3}}t_{s}^{3}(1-t_{s})^{\frac{1}{3}}}}\Bigg{]}\sin\gamma\end{split}$ (160)
So, we can rewrite the above components follow the power laws $D_{zx}^{\Omega}\propto J$ , $D_{xz}^{\Omega}\propto\frac{\alpha_{J}^{\frac{2}{3}}\mu^{(2-\frac{2}{J})}}{\hbar^{(2-\frac{2}{J})}}$ , $D_{zy}^{\Omega}\propto J$ and $D_{yz}^{\Omega}\propto\frac{\alpha_{J}^{\frac{2}{3}}\mu^{(2-\frac{2}{J})}}{\hbar^{(2-\frac{2}{J})}}$.
One can check that $D_{zx}^{m}=D_{xz}^{m}$ and $D_{zy}^{m}=D_{yz}^{m}$ and rest of the components will vanish. The contributions of all components are renormalized due to parameters $\alpha_{2}$ and $\mathit{v}_{F}$ of DWSMs as compared to single WSMs [22].
@@
Assume that an electromagnetic wave propagates in the x direction:
@@
is the Jones vector in the y-z plane with phases $\alpha_{y},\alpha_{z}$ , and the amplitudes $E_{y}=|\textbf{E}|\sin\theta$ and $E_{z}=|\textbf{E}|\cos\theta$ . Inserting Eq. (162) into Eq. (161), we obtain the nonlinear Hall current
@@
Following the same procedure of Ref.([22]), Eq.(164) can be rewritten as $j_{x}\sim(D_{yy}-D_{zz})E_{y}E_{z}$ . Hence conductivity will be
for J=1
$\sigma_{x}^{(H,B)}(2\omega)=\frac{e^{3}\tau}{(1-i\omega\tau)80\hbar^{2}}{D_{2}^{(B)}t_{s}\sin 2\gamma}$ (165)
for J=2
@@
for J=3
$\begin{split}\sigma_{x}^{(H,B)}(2\omega)&={\frac{e^{3}\tau}{\hbar^{2}(1-i\omega\tau)}}\frac{D_{2}^{(B)}}{8}\frac{\alpha_{3}^{\frac{2}{3}}\pi^{3/2}\mu^{\frac{4}{3}}}{\hbar^{\frac{4}{3}}\mathit{v_{F}^{2}}}\sin 2\gamma\biggl{[}-\frac{1}{24(3\pi)^{\frac{1}{2}}t_{s}(-1+t_{s}^{2})\Gamma{(\frac{7}{3})}\Gamma{(\frac{25}{6})}}19(1-t_{s})^{\frac{1}{3}}\Bigg{\{}5(-1+t_{s})\biggl{(}3\Bigl{(}\frac{1}{1-t_{s}}\Bigr{)}^{\frac{2}{3}}\\ &+(1-t_{s})^{\frac{1}{3}}+(1-t_{s})^{\frac{1}{3}}\biggr{)}{}_{2}F_{1}{(\frac{-2}{3},\frac{8}{3},\frac{16}{3},\frac{2t_{s}}{-1+t_{s}})}\Bigl{\{}5\biggl{(}3\Bigl{(}\frac{1}{1-t_{s}}\Bigr{)}^{\frac{2}{3}}+(1-t_{s})^{\frac{1}{3}}\biggr{)}\\ &+5(1-t_{s})^{\frac{1}{3}}t_{s}-\Bigl{(}\frac{1}{1-t_{s}}\Bigr{)}^{\frac{2}{3}}t_{s}^{2}\Bigr{\}}{}_{2}F_{1}{(\frac{1}{3},\frac{8}{3},\frac{16}{3},\frac{2t_{s}}{-1+t_{s}})}\Bigg{\}}\biggr{]}\end{split}$ (167)
Fig.(13) shows the variation of linear-B second harmonic conductivities $\sigma_{x}^{(H,B)}(2\omega)$ as function of incdent photon frequency $\omega$. In contrast to single and triple WSMs, this component has zero value for double WSM. In contrast to the case of $B=0$, the B-linear contribution to the nonlinear Hall conductivity is independent of the chirality, and the odd function of $t_{s}$ . Only in the system with broken tilt inversion symmetry ($t_{+}=t_{-}$ ), the conductivity $\sigma_{x}^{(H,B)}(2\omega)\neq 0$. From Eq. (92), evidently, $\sigma_{x}^{(H,B)}(2\omega)$ reaches its maximum when the polarization direction $\theta=\pm\pi/4$ and vanishes at $\theta=0,\pi/2$, as further reflected in Fig.(14).
Recent experiments shows the second harmonic optical response generates gaint second-order nonlinear optical polarizability, $\chi^{2}$, in transition metal monopnictides (TMMPs) such as TaAs, a class of noncentrosymmetric materials [42]. Another experment reports large non-linear Hall effect(NLHE) due to gigantic Berry curvature dipole density as generated by tilted Weyl cones near the Fermi level in a model ferroelectric Weyl semimetal in doped $(Pb_{1-x}Sn_{x})_{1-y}In_{y}Te$. The effective Berry curvature dipole derived from the experimentally observed nonlinear Hall voltage follows a scaling law with carrier concentration, which is consistent with the simplest form of the Berry curvature dipole expected for the Weyl monopoles [43].